# thanks: Equal contributionthanks: Equal contribution

### v1

@@ Results
On the other hand, we find that when ecological effects are strong ($\alpha\gg 1$), the mean times to absorption (Fig. 3a) and extinction (Fig. 3b) grow exponentially with $s$ for moderately beneficial mutants (i.e., $s\lesssim\eta$). Both of these times show a sharp peak at the critical fitness difference $s\simeq\eta/2$ (Eq. (10)). This increase in absorption time is so dramatic that it can even be directly observed in simulation trajectories of mutant and parent dynamics (see Fig. 3c–e). We further find that the maximum of both the absorption and extinction times grows exponentially with the effective population size $N_{\mathrm{eff}}$. This is in contrast with Kimura’s theory where the maximum time grows linearly with $N_{\mathrm{eff}}$. The presence of these exponentially long time scales raises the possibility that parents and mutants can transiently coexist for long periods in complex communities, highlighting a crucial difference between classical population genetics and our eco-evolutionary approach.
@@
Such coexistence is a direct consequence of frequency-dependent selection (Eq. (3)). To see this, it is helpful to think about the infinite effective population size limit of Eq. (3). In this limit, we can ignore stochasticity and focus on the fixed points of the deterministic dynamics which are given by the solutions to the equation $0=\left(s-\eta(f-f_{0})\right)f(1-f)$. When the fitness satisfies $\tilde{s}=s/\eta>1-f_{0}$ or $\tilde{s}=s/\eta<-f_{0}$, this equation has two fixed points at $f^{*}=0$ and $f^{*}=1$. On the other hand, when $f_{0}\leq\tilde{s}\leq 1-f_{0}$, this equation has an additional fixed point at $f=f_{0}+\tilde{s}\simeq\tilde{s}$, implying that the parent and mutant can potentially coexist in the absence of stochastic drift.
@@
In the presence of drift, for small and large $\tilde{s}$ — i.e., for deleterious and strongly beneficial mutants — selection almost always drives mutants to extinction (Fig. 3c) and fixation (Fig. 3e) respectively. On the other hand, for moderately beneficial mutants the presence of the additional fixed point in the deterministic limit gives rise to additional crossovers (Fig. 3g) which create a “coexistence region” between extinction and fixation, centered at frequency $f^{*}\simeq\tilde{s}$. Selection drives the mutant into this region instead of extinction or fixation, hence the parent and mutant coexist for exponentially long times (Fig. 3d) before escaping the region (SI Sec. G). Such a region is absent in classical population genetics.
@@
Parent-mutant coexistence also qualitatively explains the suppression of fixation probabilities for moderately beneficial mutants with $0<s<s_{c}$ (Fig. 2b). Recall that in classical population genetics, beneficial mutants reach fixation through establishment, i.e., crossing the drift threshold from drift to selection dominated regimes (Fig. 4a, see also Fig. 3f and h). The process of establishment is central to Kimura’s theory, since after this point, mutants grow almost deterministically towards fixation. However, this qualitative picture is modified by the coexistence region under strong ecological effects, as selection now drives mutants into this region after establishment (Fig. 3g). When the coexistence region is closer to extinction (Fig. 4b), i.e., $f^{*}<1/2$ or $s<s_{c}$, mutants that escape this region due to stochastic drift are more likely to go extinct than fix. Since such extinction occurs even after crossing the drift threshold, it suppresses fixation probabilities relative to Kimura’s prediction. On the other hand, when the coexistence region is closer to fixation (Fig. 4c), i.e., $f^{*}>1/2$ or $s>s_{c}$, it is more likely for mutants inside the region to drift to fixation instead, hence in this case coexistence does not suppress fixation.
@@ Discussion
In this study, we developed an analytic framework for integrating ecological interactions into population genetics models. Our calculations show that ecology can dramatically affect evolutionary outcomes. By combining theoretical insights from population genetics and community ecology, we show that the eco-evolutionary feedbacks give rise to an effective frequency-dependent selection between parent and mutant strains. Surprisingly, the effects of ecology are captured by a single parameter $\eta$ that can be directly calculated from ecological models using DMFT. We find that when ecological feedback is strong, the fixation probabilities for moderately beneficial mutants are suppressed compared to Kimura’s predictions (Fig. 2b). The suppression increases with effective population size (Fig. 2e), weaker parent-mutant correlations (Fig. 2e), and decreases as ecosystems become more packed (Fig. 2f). We further show that the deviations from Kimura’s formula arise from prolonged parent-mutant coexistence arising from ecological interactions (Fig. 4b and c). This coexistence results in exponentially long absorption (Fig. 3a) and extinction times (Fig. 3b) for moderately beneficial mutants, as well as coexistence regions in the phase portraits of parent-mutant dynamics (Fig. 3g).
@@
Our theory makes testable predictions for high-resolution strain tracking in natural microbiomes [59, 60]. With high enough temporal resolution, strain dynamics should exhibit intermittent stochastic fluctuations in a range of frequencies above the drift barrier (Fig. 4b–c). These dynamics should be most pronounced for moderately beneficial mutants with $s\lesssim\eta/2$. Importantly, estimating both $f^{*}$ and $s$ from observed trajectories could enable direct inference of the ecological feedback strength $\eta$ from strain dynamics alone, suggesting an interesting new empirical method for characterizing the importance of ecology in evolutionary dynamics.
@@
In the future, it will be useful to extend our theory to a wider range of ecological and evolutionary settings. On the ecological side, while our theory has so far focused on communities near steady state, ecosystems can exhibit more complex dynamical behavior such as limit cycles, multistability [61, 62], and chaos [42, 45, 63, 64, 65]. Extending our theory to these settings will enable us to understand how the fate of a mutant depends on the underlying dynamical behavior of a community. On the evolutionary side, for many species, especially microbes, mutation rates are sufficiently large so that multiple mutants can emerge before previous ones fix or go extinct [66]. Interactions between these mutants, known as clonal interference [67, 7], can significantly modify evolutionary dynamics. It will be useful to understand how such clonal interference affects the eco-evolutionary dynamics of complex ecological communities.
Materials and Methods
@@ Appendix A Lotka-Volterra model
$\left<\delta r_{i}\delta r_{j}\right>=\delta_{ij}\,,\quad\left<a_{ij}a_{kl}\right>=\delta_{ik}\delta_{jl}+\gamma\delta_{il}\delta_{jk}\,.$ (14)
The reciprocity $1\leq\gamma\leq 1$ is the correlation between the off-diagonal entries $A_{ij}$ and $A_{ji}$.
@@
We focus on the “strong-selection-weak-mutation” limit: we let the mutant invade when the community is close to a steady state, then we let the community evolve till the parent or the mutant becomes extinct. This approximation is one of the main assumptions of our analysis, which ignores the possibility of having multiple mutations in the community simultaneously.
@@
$\simeq\,$ $\sum_{ij}\int_{0}^{t}dt^{\prime}\,\nu_{ij}(t,t^{\prime})\left<(A_{mi}-A_{pi})\left(A_{jp}N_{p}(t^{\prime})+A_{jm}N_{m}(t^{\prime})\right)\right$> (28)
@@
To obtain an equation involving $f$ only, further approximations to the abundances must be made. First, we may follow the traditional convention in population genetics to restrict the total parent-mutant abundance $N_{p}+N_{m}$ to be a constant. Such restriction is further justified by the ecological steady states, in which the total abundance varies little in average, with only fluctuations from demographic noise. Therefore, we can replace
$N_{p}(t)+N_{m}(t)\simeq N_{p}(0)+N_{m}(0):=N_{0}\,$, (34)
and the effective number of individuals is $N_{\mathrm{eff}}=N_{0}/2D$. Next, since we start with the ecological steady state before the mutation, the community abundances $N_{i/pm}$ without the invasions of the parent and the mutant also vary little and we can approximate
$N_{i/pm}(t)\simeq N_{i/pm}(0)\,.$ (35)
@@
Finally, although we do not have the details of $\nu(t,t^{\prime})$, it is typically reasonable to assume the response to be very fast and can be approximated by $\nu(t,t^{\prime})\simeq\nu\delta(t-t^{\prime})$, where $\nu$ is the susceptibility of the steady state appeared in the static cavity method. With all these approximations, we arrive at
$\frac{df}{dt}=$ $\,f(1-f)\left(r_{m}-r_{p}-(1-A_{pm})N_{0}f+(1-A_{mp})N_{0}(1-f)-\sum_{i}(A_{mi}-A_{pi})N_{i/pm}(0)\right$. (36)
@@
where $s(f)$ represents frequency-dependent selection. We see that $s(f)$ is linear in $f$. To better interpret the above result, we note that $s(f)$ is in fact the difference between the growth rates between the mutant and the parent at frequency $f$. The initial difference, which is also called the invasion fitness difference, is $s(f_{0})$ and plays the same role as the selection coefficient $s$ in classic population genetics. Therefore, it is useful to rewrite $s(f)$ as a linear function in $f$ involving $s(f_{0})$:
$s(f)=s(f_{0})-\eta(f-f_{0})\,$, (39)
@@
$s(f_{0})=$ $\,r_{m}-r_{p}-(1-A_{pm})N_{0}f_{0}+(1-A_{mp})N_{0}(1-f_{0})$ (40)
$\,-\sum_{i}(A_{mi}-A_{pi})N_{i/pm}(0)+2N_{0}\gamma\sigma^{2}\nu(1-\rho)(f_{0}-1/2)\,$, (41)
@@
parametrizes the strength of the overall ecological effects. Note that the ecological feedback proportional to $\nu$ also contributes to the invasion fitness since the parent was already in the community before $t=0$. Now we can redefine $s=s(f_{0})$ following classic population genetics and get
$\frac{df}{dt}=f(1-f)\left(s-\eta(f-f_{0})\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (43)
@@
where $\phi$ is the fraction of surviving species. Since we have started with a near steady state of the community, the above expression for $\nu$ should also hold within the short time dynamics. Therefore, we also have
@@ Appendix B MacArthur consumer-resource model
$\left<\delta K_{\alpha}\delta K_{\beta}\right>=\delta_{\alpha\beta}\,,\quad\left<\delta m_{i}\delta m_{j}\right>=\delta_{ij}\,$, (53)
$\left<c_{i\alpha}c_{j\beta}\right>=\left<e_{i\alpha}e_{j\beta}\right>=\delta_{ij}\delta_{\alpha\beta}\,,\quad\left<c_{i\alpha}e_{j\beta}\right>=\kappa\delta_{ij}\delta_{\alpha\beta}\,.$ (54)
@@
$\simeq\,$ $\sum_{\alpha\beta}\int_{0}^{t}dt^{\prime}\,\chi_{\alpha\beta}(t,t^{\prime})\left<(C_{m\alpha}-C_{p\alpha})\left(E_{p\beta}N_{p}(t^{\prime})+E_{m\beta}N_{m}(t^{\prime})\right)\right$> (66)
@@
$R_{\alpha/pm}(t)\simeq R_{\alpha/pm}(0)\,,\quad N_{p}(t)+N_{m}(t)\simeq N_{0}\,,\quad\chi(t,t^{\prime})\simeq\chi\delta(t-t^{\prime})\,.$ (72)
@@
As before, we can separate the invasion fitness difference $s(f_{0})$ from $s(f)$ and rewrite the dynamics in the form of Eq. (43) as
$\frac{df}{dt}=f(1-f)\left(s-\eta(f-f_{0})\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (76)
where the invasion fitness difference is
$s=s(f_{0})=\sum_{\beta}\left(C_{m\beta}-C_{p\beta}\right)R_{\beta/pm}(0)-m_{m}+m_{p}-\eta(f_{0}-1/2)\,$, (77)

### v2

@@ Results
On the other hand, we find that when ecological effects are strong ($\alpha\gg 1$), the mean times to absorption (Fig. 3a) and extinction (Fig. 3b) become exponentially large for moderately beneficial mutants (i.e., $s_{\mathrm{inv}}\lesssim\eta$). Both of these times show a sharp peak at the critical fitness difference $s_{\mathrm{inv}}\simeq\eta/2$ (Eq. (10)). This increase in absorption time is so dramatic that it can even be directly observed in simulation trajectories of mutant and parent dynamics (see Fig. 3c–e). We further find that the maximum of both the absorption and extinction times grows exponentially with the effective population size $N_{\mathrm{eff}}$. This is in contrast with Kimura’s theory where the maximum time grows linearly with $N_{\mathrm{eff}}$. The presence of these exponentially long time scales raises the possibility that parents and mutants can transiently coexist for long periods in complex communities, highlighting a crucial difference between classical population genetics and our eco-evolutionary approach.
@@
Such coexistence is a direct consequence of frequency-dependent selection (Eq. (3)). To see this, it is helpful to think about the infinite effective population size limit of Eq. (3). In this limit, we can ignore stochasticity and focus on the fixed points of the deterministic dynamics which are given by the solutions to the equation $0=\left(s_{\mathrm{inv}}-\eta f\right)f(1-f)$. When the fitness satisfies $s_{\mathrm{inv}}>\eta$ or $s_{\mathrm{inv}}<0$, this equation has two fixed points at $f^{*}=0$ and $f^{*}=1$. On the other hand, when $0\leq s_{\mathrm{inv}}\leq\eta$, this equation has an additional fixed point at $f^{*}=\tilde{s}_{\mathrm{inv}}=s_{\mathrm{inv}}/\eta$, implying that the parent and mutant can potentially coexist in the absence of stochastic drift.
@@
In the presence of drift, for small and large $\tilde{s}_{\mathrm{inv}}$ — i.e., for deleterious and strongly beneficial mutants — selection almost always drives mutants to extinction (Fig. 3c) and fixation (Fig. 3e) respectively. On the other hand, for moderately beneficial mutants the presence of the additional fixed point in the deterministic limit gives rise to additional crossovers (Fig. 3g) which create a “coexistence region” between extinction and fixation, centered at frequency $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$. Selection drives the mutant into this region instead of extinction or fixation, hence the parent and mutant coexist for exponentially long times (Fig. 3d) before escaping the region (SI Sec. H). Such a region is absent in classical population genetics.
@@
Parent-mutant coexistence also qualitatively explains the suppression of fixation probabilities for moderately beneficial mutants with $0<s_{\mathrm{inv}}<s_{c}$ (Fig. 2b). Recall that in classical population genetics, beneficial mutants reach fixation through establishment, i.e., crossing the drift threshold from drift to selection dominated regimes (Fig. 4a, see also Fig. 3f and h). The process of establishment is central to Kimura’s theory, since after this point, mutants grow almost deterministically towards fixation. However, this qualitative picture is modified by the coexistence region under strong ecological effects, as selection now drives mutants into this region after establishment (Fig. 3g). When the coexistence region is closer to extinction (Fig. 4b), i.e., $f^{*}<1/2$ or $s_{\mathrm{inv}}<s_{c}$, mutants that escape this region due to stochastic drift are more likely to go extinct than fix. Since such extinction occurs even after crossing the drift threshold, it suppresses fixation probabilities relative to Kimura’s prediction. On the other hand, when the coexistence region is closer to fixation (Fig. 4c), i.e., $f^{*}>1/2$ or $s_{\mathrm{inv}}>s_{c}$, it is more likely for mutants inside the region to drift to fixation instead, hence in this case coexistence does not suppress fixation.
@@ Discussion
In this study, we developed an analytic framework for integrating ecological interactions with a complex community into population genetics models. Our calculations show that interactions with a highly-diverse community can dramatically affect evolutionary outcomes. Despite their complexity, ecological feedbacks give rise to a remarkably simple form of frequency-dependent selection between parent and mutant strains. Surprisingly, the entire community’s frequency-dependence can be captured by a single parameter $\eta$ which can be easily calculated from the community for a wide variety of ecological models. Thus our results reveal a principled method to apply existing population genetics models with frequency-dependent selection even in the presence of complex communities.
We find that when ecological feedback is strong, the fixation probabilities for moderately beneficial mutants are suppressed compared to Kimura’s predictions (Fig. 2b). The suppression increases with effective population size (Fig. 2e), weaker parent-mutant correlations (Fig. 2e), and decreases as ecosystems become more packed (Fig. 2f). We further show that the deviations from Kimura’s formula arise from prolonged parent-mutant coexistence arising from ecological interactions (Fig. 4b and c). This coexistence results in exponentially long absorption (Fig. 3a) and extinction times (Fig. 3b) for moderately beneficial mutants, as well as coexistence regions in the phase portraits of parent-mutant dynamics (Fig. 3g).
@@
Our theory makes testable predictions for high-resolution strain tracking in natural microbiomes [65, 66]. With high enough temporal resolution, strain dynamics should exhibit intermittent stochastic fluctuations in a range of frequencies above the drift barrier (Fig. 4b–c). These dynamics should be most pronounced for moderately beneficial mutants with $s_{\mathrm{inv}}\lesssim\eta/2$. Importantly, estimating both $f^{*}$ and $s_{\mathrm{inv}}$ from observed trajectories could enable direct inference of the ecological feedback strength $\eta$ from strain dynamics alone, suggesting an interesting new empirical method to quantify the impact of ecology during evolutionary dynamics [67].
@@
In the future, it will be useful to extend our theory to a wider range of ecological and evolutionary settings. On the ecological side, while our theory has so far focused on communities near steady state, ecosystems can exhibit more complex dynamical behavior such as limit cycles, multistability [68, 69], and chaos [42, 45, 70, 71, 72]. Extending our theory to these settings will enable us to understand how the fate of a mutant depends on the underlying dynamical behavior of a community. On the evolutionary side, for many species, especially microbes, mutation rates are sufficiently large so that multiple mutants can emerge before previous ones fix or go extinct [73]. In particular, new mutants can easily emerge when previous mutants are still coexisting transiently with their parents. Interactions between these mutants, known as clonal interference [74, 7], can significantly modify evolutionary dynamics. It will be useful to understand how such clonal interference affects the eco-evolutionary dynamics of complex ecological communities. Further, over longer timescales eco-evolutionary feedbacks may result in communities with structured ecological interactions compared with the randomly assembled communities we have focused on in this paper [14, 75, 76, 15, 39, 48]. It will also be useful to understand how to generalize our results to communities evolved over the long-term.
Methods
@@ Appendix A Lotka-Volterra model
$\left[\delta r_{i}\delta r_{j}\right]=\delta_{ij}\,,\quad\left[a_{ij}a_{kl}\right]=\delta_{ik}\delta_{jl}+\gamma\delta_{il}\delta_{jk}\,$, (14)
where the square brackets denote averages over random parameters instead of the demographic noise. The reciprocity $1\leq\gamma\leq 1$ is the correlation between the off-diagonal entries $A_{ij}$ and $A_{ji}$. The results below also apply to other distributions with similarly defined mean and variance.
@@
We focus on the “strong-selection-weak-mutation” limit: we assume an ecological steady state and let the mutant invade when the community is close to the steady state, then we let the community evolve till the parent or the mutant becomes extinct, even though the parent and mutant may transiently coexist for a long time. This approximation is one of the main assumptions of our analysis, which ignores the possibility of having multiple mutations in the community simultaneously due to finite mutation rates.
@@
$\simeq\,$ $\sum_{ij}\int_{0}^{t}dt^{\prime}\,\nu_{ij}(t,t^{\prime})\left[(A_{mi}-A_{pi})\left(A_{jp}N_{p}(t^{\prime})+A_{jm}N_{m}(t^{\prime})\right)\right]$ (28)
@@
To obtain an equation involving $f$ only, further approximations to the abundances must be made. First, we may follow the traditional convention in population genetics to restrict the total parent-mutant abundance $N_{p}+N_{m}$ to be a constant. Such restriction is further justified when the parent-mutant correlation $\rho$ is high. As discussed in Sec. F, the total parent-mutant abundance varies little in average when $\rho$ is high and the whole community is near the ecological steady state. There are only fluctuations from demographic noise. Therefore, we can replace
$N_{p}(t)+N_{m}(t)\simeq\left<N_{p}(0)+N_{m}(0)\right>:=N_{0}\,$, (34)
and the effective number of individuals is $N_{\mathrm{eff}}=N_{0}/2D$. We recall that the angle bracket denotes the average over demographic fluctuations in the initial ecological steady state.
Next, since we start with the ecological steady state before the mutation, the community abundances $N_{i/pm}$ without the invasions of the parent and the mutant also vary little. These variations are further averaged out when $N_{i/pm}$ enters the sum in Eq. (32). Similar to the above, we can approximate
$\sum_{i}(A_{mi}-A_{pi})N_{i/pm}(t)\simeq\sum_{i}(A_{mi}-A_{pi})\left<N_{i/pm}(0)\right$>\,. (35)
@@
Finally, although we do not have the full quantitative details of $\nu(t,t^{\prime})$, it is typically reasonable to assume the response to be much faster than the timescale of growth rates [41], especially when the community is near the ecological steady state. Empirically, we observe that to predict mutation outcomes, it suffices to approximate the susceptibility simply by $\nu(t,t^{\prime})\simeq\nu\delta(t-t^{\prime})$, where $\nu$ is the susceptibility of the steady state appeared in the static cavity method. As we will see, the approximation becomes less accurate when there are more indirect interactions in the ecological dynamics (Sec. E) or the parent-mutant correlation is not high (Sec. F). Quantifying the regimes where this approximation is controllable is left for future work.
With all these approximations, we arrive at
[…]
@@
where $s(f)$ represents frequency-dependent selection. We see that $s(f)$ is linear in $f$. To better interpret the above result, we note that $s(f)$ is in fact the difference between the growth rates between the mutant and the parent at frequency $f$. The difference $s(f=0)$ is also the invasion fitness of the mutant $s_{\mathrm{inv}}$ and plays the same role as the selection coefficient $s$ in classic population genetics. Therefore, it is useful to rewrite $s(f)$ as a linear function in $f$ involving $s_{\mathrm{inv}}$:
$s(f)=s_{\mathrm{inv}}-\eta f\,$, (39)
@@
$s_{\mathrm{inv}}=r_{m}-r_{p}+(1-A_{mp})N_{0}-\sum_{i}(A_{mi}-A_{pi})\left<N_{i/pm}(0)\right>-N_{0}\gamma\sigma^{2}\nu(1-\rho)\,$, (40)
@@
parametrizes the strength of the overall ecological effects. Note that the ecological feedback proportional to $\nu$ also contributes to the invasion fitness since the parent was already in the community before $t=0$. In conclusion, we get
$\frac{df}{dt}=f(1-f)\left(s_{\mathrm{inv}}-\eta f\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (42)
@@
where $\phi$ is the fraction of surviving species. Since we have started with a near steady state of the community for the mutation, the above expression for $\nu$ should also hold approximately after the mutation. Therefore, we also have
@@ Appendix B MacArthur consumer-resource model
$\left[\delta K_{\alpha}\delta K_{\beta}\right]=\delta_{\alpha\beta}\,,\quad\left[\delta m_{i}\delta m_{j}\right]=\delta_{ij}\,$, (52)
$\left[c_{i\alpha}c_{j\beta}\right]=\left[e_{i\alpha}e_{j\beta}\right]=\delta_{ij}\delta_{\alpha\beta}\,,\quad\left[c_{i\alpha}e_{j\beta}\right]=\kappa\delta_{ij}\delta_{\alpha\beta}\,.$ (53)
@@
$\simeq\,$ $\sum_{\alpha\beta}\int_{0}^{t}dt^{\prime}\,\chi_{\alpha\beta}(t,t^{\prime})\left[(C_{m\alpha}-C_{p\alpha})\left(E_{p\beta}N_{p}(t^{\prime})+E_{m\beta}N_{m}(t^{\prime})\right)\right]$ (65)
@@
$\sum_{\beta}(C_{m\beta}-C_{p\beta})R_{\alpha/pm}(t)\simeq\sum_{\beta}(C_{m\beta}-C_{p\beta})\left<R_{\alpha/pm}(0)\right>\,,\quad N_{p}(t)+N_{m}(t)\simeq N_{0}\,,\quad\chi(t,t^{\prime})\simeq\chi\delta(t-t^{\prime})\,.$ (71)
@@
As before, we can separate the invasion fitness $s_{\mathrm{inv}}=s(f=0)$ from $s(f)$ and rewrite the dynamics in the form of Eq. (42) as
$\frac{df}{dt}=f(1-f)\left(s_{\mathrm{inv}}-\eta f\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (75)
where the invasion fitness is
$s_{\mathrm{inv}}=\sum_{\beta}\left(C_{m\beta}-C_{p\beta}\right)\left<R_{\beta/pm}(0)\right>-m_{m}+m_{p}+\frac{\eta}{2}\,$, (76)